OpenStudy (anonymous):

Abdul is pushing his 20.0 kg little brother on a merry-go-round with a radius of 2.0 meters. If his little brother is on the edge of the merry-go-round and has an angular momentum of 360 kg m2/s, what is his speed?

OpenStudy (anonymous):

last one! @shamim

OpenStudy (shamim):

Given m=20kg r=2 L=360kg.m/s v=?

OpenStudy (shamim):

Use L=mvr 360=20*v*2 v=?

OpenStudy (anonymous):

i'm smart